# Studying the Effects of Collaboration in Interactive Theme Discovery Systems

### v1

@@ 7 Limitations
The study presented in this paper has three main limitations.
(1) While we selected two distinct tools to perform our analysis of synchronous vs. asynchronous settings, these tools are not representative of the full range of solutions that have been proposed to support qualitative coding. In particular, neither of these systems utilize large language models (LLMs), which have been at the forefront of recent developments in the field Dai et al. (2023); Gao et al. (2024), and which have garnered significant interest from qualitative researchers Hayes (2023); Roberts et al. (2024); Aubin Le Quéré et al. (2024).
(2) As prior work has shown Hoyle et al. (2021), automated metrics and human evaluation do not always correlate well for unsupervised or semi-supervised theme induction tasks like topic modeling. Although we did incorporate two types of human evaluation: a post-hoc manual verification (Sec 5.3) and a user study (Sec 5.4), both of these evaluations were limited in scope. This is primarily due to how expensive and time consuming it is to perform an exhaustive human evaluation. While we are confident in the findings of our study, future work should expand on the human analysis of these solutions to provide a more robust view.
@@ 8 Ethical Considerations
All information needed to replicate our experiments is presented in the paper. We reported all experimental settings, as well as any pre-processing steps, learning configurations, hyper-parameters, and additional technical details. Due to space constraints, some of this information was relegated to the Appendix. In addition to this, we will make the results of the annotation experiment available to the community, as well as the code to produce all of our reported results***https://github.com/blast-cu/interactive-systems .
Appendix A Interactive System Details
@@
Finding Partitions Experts can find partitions in the space of unassigned instances. We currently support the K-means Jin2010 and Hierarchical Density-Based Clustering mcinnes2017hdbscan algorithms.
@@
Visualizing Global Explanations Experts can visualize aggregated statistics and explanations for the global state of the system. To do this, we aggregate all instances in the database. Explanations include theme distribution, coverage statistics, and t-sne plots vanDerMaaten2008.

### v2

@@ 8 Limitations
The study presented in this paper has four main limitations.
(1) While we selected three distinct, representative tools to perform our analysis of synchronous vs. asynchronous settings, the list is of course non-exhaustive. A larger study incorporating more tools could yield additional insights.
(2) As prior work has shown Hoyle et al. (2021), automated metrics and human evaluation do not always correlate well for unsupervised or semi-supervised theme induction tasks like topic modeling. Although we did incorporate two types of human evaluation: a post-hoc manual verification (Sec 5) and a user study (Sec 6), both of these evaluations were limited in scope. This is primarily due to how expensive and time consuming it is to perform an exhaustive human evaluation. While we are confident in the findings of our study, a larger study could further strengthen our conclusions.
(3) We used a single dataset in our experiments, which is composed of short English texts. While we believe that this was sufficient to arrive at strong conclusions across systems and collaboration modalities, datasets with different characteristics (e.g., longer documents, non-English language) could pose additional challenges.
@@ 9 Ethical Considerations
All information needed to replicate our experiments is presented in the paper. We reported all experimental settings, as well as any pre-processing steps, learning configurations, hyper-parameters, and additional technical details. Due to space constraints, some of this information was relegated to the Appendix. In addition to this, we will make the results of the annotation experiment available to the community, as well as the code to produce all of our reported results..
Appendix A Topic Model Experimental Settings
To ensure comparable results, we use the same starting distribution of 13 topics for all our experiments using the same hyperparameters as Fang et al. (2021), which are $\alpha=1.0$, $\beta=0.5$, $\gamma=1.5$. The QDTM also allows queries to be input prior to topic model initialization to produce partitions that follow prior knowledge, but we do not take advantage of this capability. The same initial topic model is provided for all experiments.
Appendix B Experimental Settings and Interactive System Details
During the interactive coding process, researchers are provided initial clusters to identify themes. We use $K=10$ means clustering to generate initial partitions and the same partitions are provided for all experiments. Users select positive and negative examples for each theme, which are used to create a distributional representations to calculate semantic similarity with unlabeled documents. Simultaneously, supporting concepts are defined using logical rules which researchers can use to label example documents. The relationship between themes and concepts are then used as a structured inference task to predict label assignments. By using multiple sources of information, the model can much more efficiently extrapolate from a small set of manually labeled data.
B.1 Relational Approach Operational Details
@@
Finding Partitions Experts can find partitions in the space of unassigned instances. We currently support the K-means Jin and Han (2010) and Hierarchical Density-Based Clustering Mendelsohn et al. (2021) algorithms.
@@
Visualizing Global Explanations Experts can visualize aggregated statistics and explanations for the global state of the system. To do this, we aggregate all instances in the database. Explanations include theme distribution, coverage statistics, and t-sne plots van der Maaten and Hinton (2008).
@@
Appendix C Semi-Structured Interview
C.1 Interviewing
We usually administered interviews after annotation sessions. In some cases, due to time constraints from the annotation session taking longer than expected, we were unable to pose these questions to the annotators. Asynchronous annotators were asked questions individually about their experience, whereas synchronous annotator groups were usually asked questions with their fellow annotators.
C.2 Script
  1. 1.
    Have you worked on annotation projects before? Did these annotation projects use qualitative coding strategies (ex: grounded theory)? How experienced are you as an annotator?
  2. 2.
    How was your experience on the COVID-19 vaccine annotation session we conducted on Sunday? Particularly, we are interested in your thoughts and feelings over the session.
  3. 3.
    You annotated in a group, working together as a team. Did you find this setup to be beneficial? What were some of the limitations you faced, both individually and and as a group, when working synchronously?
  4. 4.
    On a similar line, what would you consider to be the pros and cons if you were to annotate alone?
The last questions would be flipped based on if we are posing it to synchronous or asynchronous annotators.
Appendix D LLM-based Experimental Configuration and Prompt Details
For this study, we use the same starting partitions as used in the relational approach experiment in App. B. The original work uses existing codes from a theoretical framework whereas we use codes defined by the user during this step, but this does not affect the overall process.
code-book Gwet’s $AC_{1}$ Unlabeled Docs
Sync 1 $0.42$ $5,548(6.5\%)$
Sync 2 $0.49$ $9,766(11.4\%)$
Async 1 $0.61$ $611(0.7\%)$
Async 2 $0.62$ $2,506(2.9\%)$
Async 3 $0.46$ $13,816(16.1\%)$
Table 4: Results for the selected prompt for each coding session using the LLM-based system. Gwet’s $AC_{1}$ is used to select the best prompt for running the full dataset. The number of unlabeled documents represent documents where the LLM produced a label not created by human annotators after running the full dataset (percentage of the dataset unlabeled).
Using the LLM, we generate 3 additional templates based on a prompt structure provided in the original work. We used the Llama 3.2 3B-Instruct (Grattafiori et al., 2024) model for all generation tasks:
D.1 LLM Hyperparameters
Batch size: 32
Model: Llama 3.2 3B-Instruct
GPU: A100 40GB VRAM
Average Compute Time: 24hrs per job
Number of jobs: 5
D.2 LLM Prompts
To code this tweet, do the following:
- First, read the codebook and the tweet.
- Next, decide which code is most applicable and explain your reasoning for the coding decision.
- Finally, generate json with your code and your reason for the coding decision. The respone MUST be formatted as JSON.
Codes:
<codes>
Codebook:
<codebook>
Tweet:
<tweet>
JSON Output:
"code" : "",
"reason" : ""
To code this tweet, do the following:
First, read the codebook and the tweet.
Next, decide which code is most applicable based on the tweet’s content and explain your reasoning for the coding decision.
Finally, generate a JSON object with the selected code and provide a brief explanation for your coding decision.
The response MUST be formatted as JSON.
Codebook: Themes: <"theme": "definition">
Tweet: < "text": "<text>" >
JSON Output: < "code": "", "reason": "" >
To generate code for this tweet, provide a step-by-step explanation of how to approach the task.
First, analyze the tweet’s content and identify key concepts, such as the type of object or class being described, any specific behaviors or requirements, and relevant keywords.
Next, evaluate the codebook options and determine which one is most applicable. Explain your reasoning for your decision, including any similarities between the tweet and the code definitions, or any specific requirements mentioned in the tweet that align with a particular code.
Finally, generate a JSON object with the selected code and provide additional context, including:
A clear explanation of how you arrived at your chosen code
Any relevant notes or comments about the code’s functionality and requirements
A brief comparison to other codes in the book, if applicable
The response MUST be formatted as JSON.
Codebook: <codebook>
Tweet: <tweet>
JSON Output: < "code": "", "reasoning": "", "context": "" >
To analyze this tweet and select a relevant theme, follow these steps:
First, read the tweet and identify key concepts, such as emotions, objects, or ideas mentioned in the text.
Next, evaluate the theme options and determine which one is most applicable. Explain your reasoning for your decision, including any connections you see between the tweet’s content and the theme definitions.
Then, generate a JSON object with the selected theme and provide additional insight into your analysis. Include:
A clear explanation of how you arrived at your chosen theme
Any specific characteristics or keywords from the tweet that support your decision
A brief comparison to other themes, if applicable
The response MUST be formatted as JSON.
Themes: <Codebook>
Tweet: <tweet>
JSON Output: <"theme": "", "insight": "">